lilbunnybun:

math 9th grade ss💔🌹😖

localcheerleader:

ss pls (:

localcheerleader:

so number 1 right?

lilbunnybun:

Yes

localcheerleader:

get y on a side all by itself y+2=5

lilbunnybun:

𝑦 + 2 − 2 = 5 − 2

localcheerleader:

and we want y by self y+2=5 -2=-2


y =3

localcheerleader:

yes (:

lilbunnybun:

𝑦 = 3

lilbunnybun:

Ok

localcheerleader:

yep

localcheerleader:

need more ask

lilbunnybun:

3 + 𝑥 = − 5 {\color{#c92786}{3+x}}=-5 3+x=−5 𝑥 + 3 = − 5

lilbunnybun:

𝑥 = − 8

localcheerleader:

so you minus 3 from both sides- yes

lilbunnybun:

k

lilbunnybun:

ok so i got those down

lilbunnybun:

can u give me another medal-

localcheerleader:

okay and yw

lilbunnybun:

thx

localcheerleader:

yw